# Unc-TTP: A Method for Classifying LLM Uncertainty to Improve In-Context Example Selection

### v2

@@ 5.1 Experiment Settings
As for the uncertainty category, we enumerate all possible combinations of outcomes in the three samplings, with at most six for the Unc-TTP method. We choose the one category with the highest accuracy on the validation set as the final in-context examples for testing. It should be noted that our approach is a one-example-for-all method. Once the selection of the uncertainty category is made, the same example is applied to all test instances, eliminating the need to retrieve the most compatible one for each test instance as in previous approaches Robertson and Zaragoza (2009); Liu et al. (2021); Nguyen et al. (2022); Qin et al. (2023).
Implementation Details
The dataset and model configurations used in this part of the experiment are consistent with the setting from Section 4.1. For the sampling process, we repeated three times for each category using different random seeds and reported the average accuracy along with standard deviation. We instruct LLMs with the same prompt as in the no-label setting to evaluate their performance under demonstrations.
Comparison Methods
@@
We evaluate our approach in comparison with the following methods for selecting ICL demonstrations: Random randomly selects examples from the training set as ICL demonstrations. BM25 Robertson and Zaragoza (2009) ranks the training examples based on their relevance to each test instance. The top-ranked examples are then selected as ICL demonstrations. Similarity  Liu et al. (2021) selects the top-$k$ semantically similar examples from the training set as demonstrations for each test sample. Following Qin et al. (2023), all samples are encoded by Sentence-BERT Reimers and Gurevych (2019) to obtain representations for calculating the cosine similarity.
SelfCheckGPT Manakul et al. (2023a), a sampling-based method that applies in the fact-checking task. It uses a rhetorical question on the corresponding answers after the sampling to verify whether the model supports the answer and scores each instance by the consistency of verification. The higher the score, the more likely the model’s answer is hallucinated; otherwise, it is factual. We use the score computed from the three samplings to guide the in-context example selection.
Sampling-based method is inspired by self-consistency  Wang et al. (2023b) and employs $q$ decoding paths, followed by majority voting to identify the instances that the LLM is most certain of. We extend this method by classifying the sampling results into CertainR, CertainW, and Uncertain. The Uncertain includes two categories which are correct once or twice out of $q=3$ samplings.
5.2 Main Results: 1-Shot ICL
Methods Cat. SH HS FP Average
@@ 5.2 Main Results: 1-Shot ICL
Random - 69.8 (1.6) 53.5 (1.5) 78.3 (1.8) 67.2 (1.6)
Retrieval BM25 68.0 (0.7) 55.4 (1.1) 74.7 (1.2) 66.0 (1.0)
Sim 69.9 (0.7) 52.7 (1.5) 77.3 (0.0) 66.6 (0.7)
SelfCheck MaxS 66.2 (1.7) 57.3 (1.9) 77.8 (1.1) 67.1 (1.6)
GPT MinS 66.0 (2.5) 56.7 (1.2) 77.3 (1.3) 66.7 (1.7)
Sampling CerW 67.9 (2.7) 53.3 (2.1) 78.8 (0.2) 66.7 (1.7)
CerR 66.9 (2.4) 53.8 (1.8) 80.0 (1.9) 66.9 (2.0)
Unc 69.1 (1.0) 53.8 (1.0) 79.8 (1.2) 67.6 (1.1)
Unc-TTP (ours) CerW 68.5 (2.0) 59.5 (9.5) 79.1 (1.7) 69.0 (4.4)
CerR 70.4 (1.9) 56.9 (2.9) 77.0 (3.4) 68.1 (2.7)
Unc 71.2 (2.1) 58.8 (3.2) 80.8 (1.3) 70.3 (2.2)
@@
Random - 71.2 (1.6) 81.6 (4.4) 84.6 (2.3) 79.1 (2.8)
Retrieval BM25 72.8 (0.6) 80.0 (1.1) 85.5 (0.9) 79.4 (0.9)
Sim 76.0 (0.5) 78.8 (0.6) 86.1 (0.3) 80.3 (0.5)
SelfCheck MaxS 65.5 (2.9) 80.8 (2.1) 82.9 (1.5) 76.4 (2.2)
GPT MinS 70.2 (3.0) 82.0 (1.4) 80.2 (2.9) 77.5 (2.4)
Sampling CerW 65.7 (5.1) 81.3 (1.8) 87.3 (1.9) 78.1 (2.9)
CerR 71.2 (3.8) 77.3 (5.0) 86.0 (1.4) 78.2 (3.4)
Unc 72.0 (0.9) 79.5 (1.4) 87.0 (1.2) 79.5 (1.2)
Unc-TTP (ours) CerW 71.2 (5.4) 81.0 (2.3) 86.4 (2.4) 79.5 (3.4)
CerR 66.6 (3.8) 74.6 (2.0) 84.3 (0.9) 75.2 (2.2)
Unc 74.7 (0.8) 81.8 (2.9) 88.1 (0.8) 81.5 (1.5)
GPT-3.5
Random - 83.1 (4.3) 88.8 (1.8) 83.2 (0.8) 85.0 (2.3)
Retrieval BM25 74.8 (1.6) 87.2 (1.0) 83.2 (1.0) 81.7 (1.2)
Sim 79.2 (0.6) 87.3 (0.6) 83.4 (0.2) 83.3 (0.5)
  SelfCheck MaxS 77.3 (3.5) 87.5 (0.7) 81.0 (0.7) 81.9 (1.6)
GPT MinS 78.7 (2.9) 89.8 (0.5) 82.9 (0.6) 83.8 (1.3)
Sampling CerW 78.5 (3.9) 91.7 (0.8) 83.0 (1.1) 84.4 (1.9)
CerR 77.3 (2.9) 91.0 (1.3) 82.7 (1.8) 83.7 (2.0)
Unc 80.1 (5.5) 92.5 (1.8) 83.5 (1.3) 85.4 (2.9)
Unc-TTP (ours) CerW 78.9 (3.6) 88.2 (1.8) 79.8 (1.4) 82.3 (2.3)
CerR 82.8 (2.9) 88.8 (2.4) 82.1 (2.0) 84.6 (2.4)
Unc 77.8 (2.4) 93.0 (0.5) 84.8 (1.7) 85.2 (1.5)
Table 1: 1-shot results of LLMs on three subjective tasks in the test set. The table presents the mean accuracy across three estimations, with standard deviations in parentheses. The highest accuracy values under each method are underlined. The best results for each dataset are highlighted in bold. Note that Retrieval* refers to methods that select different examples for each test instance. In Retrieval, Sim denotes Similarity. MaxS and MinS represent self-check scores. Cer denotes Certain, while Unc denotes Uncertain. Cat. refers to Category.
The test results for 1-shot experiments on three datasets are shown in Table 1. According to an analysis of the uncertainty category used for Unc-TTP and the sampling-based method, as described in Section 4.2, 010 and 011 are the most informative categories for small models including Llama-2 and Mistral.
The results demonstrate that the Uncertainty category of Unc-TTP and the sampling-based method outperform their Certain ones (mean of CertainR and CertainW) by an average of 2.55 points and 1.16 points across three models and datasets. This highlights the significant informativeness of uncertain instances when used as in-context examples.
Uncertainty-based sampling methods also show a clear advantage over other baselines. Compared to retrieval-based methods like BM25 and Similarity, Unc-TTP achieves superior performance with an average improvement of 3.27 points and 2.26 points, respectively, while sampling-based methods show gains of 1.74 points and 0.73 points. Against a random baseline, Unc-TTP surpasses it by 1.88 points, whereas the sampling-based method exceeds it by only 0.36 points. These findings suggest that selecting in-context examples with higher uncertainty can lead to a notable enhancement in LLMs’ performance.
Although both uncertainty sampling methods demonstrate significant advantages, Unc-TTP, with its stricter certainty classification, delivers superior and more robust results compared to the sampling-based method. Specifically, the accuracy of the Unc-TTP-guided uncertainty sampling method surpasses that of the sampling-based method by an average of 1.52 points on three models and datasets.
As for SelfCheckGPT, the overall results are not as robust as those of Unc-TTP and the sampling-based method. Instances with high or low scores do not exhibit clear advantages when used as in-context examples. This suggests that instances selected through self-checking may not provide substantial informativeness for ICL.
Despite the overall success of our approach, the accuracy of GPT-3.5 on the SH dataset is modest. One possible reason is that, compared to other LLMs or datasets, GPT-3.5 only has two types of uncertainty examples in SH, so the category selected based on the validation set is likely not optimal.
5.3 N-Shot Results
We conduct $N$-shot experiments with Mistral to evaluate whether Unc-TTP-guided uncertainty sampling can maintain its superiority as the number of shots increases. The results, depicted in Figure 5, reveal that uncertain examples identified by Unc-TTP are generally more informative than those selected by the sampling-based method or chosen randomly. Unc-TTP consistently outperforms other baselines across 1 to 4 shots, highlighting the informativeness of the examples selected under Unc-TTP’s guidance. However, the benefits of uncertainty sampling are potentially marginal. The advantages of uncertainty sampling diminish for the SH and FP datasets when the number of shots exceeds 8. We argue that the marginal benefits of uncertainty sampling vary depending on the dataset and the length of the selected instances. The advantages of uncertain examples identified by Unc-TTP are evident in the smaller number of shots.
5.4 Discussion
More capable LLMs have demonstrated a stronger ability to resist skepticism or incorrect information provided by users, and they maintain their stance while achieving a higher accuracy rate. Therefore, we seek to investigate whether the uncertain instances where stronger LLMs waver are more informative and thus provide stronger guidance for weaker LLMs compared to the self-guided ones.
We implement Unc-TTP with GPT-4 and leverage the examples selected under the guidance of the uncertainty of GPT-4 as the demonstrations for Llama-2 and Mistral. The results, as presented in Table 2, reveal that although the GPT-4’s uncertainty-guided example selection occasionally outperforms the self-guided selection of weaker models, the advantage is inconsistent. In contrast, the accuracy of weak models with self-guided uncertainty sampling is higher in 5 out of 6 cases and is more robust than the random sampling baseline. This indicates that relying on a stronger model’s uncertainty is unnecessary for in-context example selection.
Type SH HS FP
Mistral
GPT4-Unc 69.4 (2.7) 81.8 (1.8) 89.8 (3.3)
Self-Unc 69.6 (2.0) 82.0 (2.6) 91.1 (1.0)
Random 67.3 (1.2) 81.8 (3.2) 89.2 (4.3)
Llama-2
GPT4-Unc 67.0 (1.7) 58.4 (1.4) 87.7 (1.5)
Self-Unc 68.0 (1.2) 62.4 (1.8) 86.5 (1.9)
Random 66.8 (1.7) 53.4 (3.3) 84.4 (3.2)
Table 2: Comparative results of Unc-TTP-guided uncertainty sampling on ICL: uncertainty of GPT-4 vs. self-guided uncertainty of Llama-2 and Mistral.
@@ 6 Conclusion
We propose a novel uncertainty testing paradigm, Unc-TTP, for classifying the uncertainty levels of both open- and closed-source LLMs. This paradigm involves enumerating all possible LLM output combinations under three testing scenarios and classifying uncertainty levels based on output consistency, thus exposing areas where LLM are not certain and potentially hallucinated. We provide a detailed analysis of uncertainty properties classified by Unc-TTP and the sampling-based method, noting that more capable models exhibit fewer proportions of wavering and greater certainty. Moreover, we introduce a self-guided, in-context example selection strategy that utilizes the uncertainty obtained by Unc-TTP. Our approach selects more informative examples than the sampling-based and retrieval-based approaches, achieving excellent results across all three models. Our work has shed light on the estimation of LLM uncertainty and laid the ground for utilizing the extrinsic uncertainty manifestation of LLM to improve ICL performance.
@@
Firstly, our proposed Unc-TTP inherits the prompt-sensitive nature of the LLMs, which may introduce risks when applying Unc-TTP-guided uncertainty sampling.
Secondly, we have only evaluated Unc-TTP on subjective tasks. Given highly subjective tasks often require a deep understanding of contextual subtleties, their labels are sometimes "biased" themselves depending on the annotator decomposition (Li et al., 2022, 2023b). This characteristic is more compatible with Unc-TTP, as it can effectively expose areas where LLMs exhibit uncertainty compared to objective reasoning tasks. For the implementation of Unc-TTP on subjective tasks will be a primary focus in our future work.
Thirdly, we have only tested Unc-TTP on classification tasks due to the limited time. However, it should be noted that our approach is not limited to the categorization task. On the generative task, we can similarly transform the model’s answer into a binary fact-checking task and obtain the tripartite paradigm by means of different interrogative formulations. This will be our future work.
Finally, due to limitations in time and resources, we have not yet tested the combination of uncertainty-based examples and randomly selected examples in K-way N-shot ICL settings. This aspect will also be addressed in our future work.
@@
Appendix A Dataset Implementation Details
For the FP dataset, each sentence is annotated between 5 to 8 times. To ensure high data quality, we included only those sentences with 100% annotation agreement, resulting in 2,264 examples available for training and evaluation. Among these, only 303 examples were labeled as negative. Consequently, we randomly selected 100 instances for each label to form the training set and 200 instances for each label for the validation set. We randomly sample 200 instances from the split that reach 75% annotation agreement for the test set.
  • SH: sarcastic / non-sarcastic
  • HS: humorous / not humorous
  • FP: positive / neutral / negative
  Dataset Train Valid Test Label
SH 500 1500 200 2
HS 500 1500 200 2
FP 300 600 300 3
Appendix B Detailed Results of the LLM Uncertainty Assessment
In this section, we present the data distribution of Unc-TTP on the FP dataset in Table 5.

### v4

@@ 3.3 Example Selection with Unc-TTP
4 Experimamtal Setup
Figure 4: The relative improvement in accuracy of inconsistency-defined uncertainty examples compared to certainty examples in ICL for each model. The uncertainty examples from sampling-based methods (the bottom 2 rows) demonstrate greater robustness than those from verification-based methods (the top 2 rows), where positive accuracy gains (highlighted in red) are observed more consistently than negative ones (highlighted in blue).
4.1 Model Configurations
We perform our experiments with GPT-3.5 and GPT-4  [^1] Our experiments were done with gpt-3.5-turbo-1106 and the gpt-4-0613 checkpoint. The results for certain examples may change due to model updates.  (OpenAI, 2024), llama-2-7b-chat (Touvron et al., 2023) and mistral-7b-instruct (Jiang et al., 2023), a model that significantly outperforms the llama-2-13b-chat  (Jiang et al., 2023). All generations are done under the temperature of 0.7 except Unc-TTP example selection is done via greedy decoding.
4.2 Datasets
We select three representative subjective text classification tasks for our experiments, according to the subjectivity ranking Li et al. (2023). Sarcasm Headlines (SH) (Misra and Arora, 2023), a high-quality news headlines sarcasm detection dataset. Humor Speech (HS) (Annamoradnejad and Zoghi, 2020), a dataset consisting of short texts with both humorous and non-humorous content. Financial Phrasebank (FP) (Malo et al., 2014), a sentiment classification task that divides finance-related sentences as positive, negative, or neutral. Please refer to Appendix B.1 for a detailed data description.
4.3 Comparing Example Selection Methods
4.3.1 Output Inconsistency-based Methods
To evaluate the validity of inconsistency-defined uncertainty and compare the informativeness of uncertainty versus certainty instances, we implement two types of inconsistency measurement methods: Verification-based and Sampling-based methods, with $q=3$ to align with the Unc-TTP tripartite testing paradigm.
Verification-based Methods
These two-stage methods measure consistency between the LLM’s answer and its self-verification, checking if the model consistently supports its initial response. We adopt SelfCheckGPT Manakul et al. (2023) and P(True) Kadavath et al. (2022), both of which are usually used in fact-checking tasks. SelfCheckGPT scores each instance based on the consistency across $q$ rounds of verification. The higher scores indicate greater uncertainty (denoted as UncS), while lower scores reflect higher certainty (denoted as CerS). P(True) calculates the probability that the LLM outputs "True" during the verification stage, where lower probabilities indicate uncertainty, and higher probabilities suggest certainty.
Sampling-based Methods
These assess output inconsistency through multiple sampling iterations, with Unc-TTP falling into this category. The Vanilla Sampling (Vanilla) method, also known as temperature sampling, inspired by self-consistency Wang et al. (2023b), uses $q$ decoding paths followed by majority voting to identify instances of highest certainty. We extend this approach by classifying output from the $q$ samplings into the categories of CerR, CerW, and Unc.
4.3.2 Traditional Methods
To evaluate the effect of our proposed Unc-TTP-guided uncertainty as an active in-context example selection strategy, we compare our method with the Random baseline as well as other common AL approaches, following Margatina et al. (2023).
@@ 4.3.2 Traditional Methods
Similarity Liu et al. (2021) selects the top-$k$ semantically similar examples as demonstrations for each test sample. BM25 Robertson and Zaragoza (2009) ranks training examples based on their relevance to each test instance, and the top-ranked examples are selected as ICL demonstrations. Diversity Margatina et al. (2023) encodes all train data with Sentence-BERT Reimers and Gurevych (2019) and performs $k$-means clustering. We select one instance from each cluster. Perplexity and Entropy are two classical measures to evaluate LLMs’ uncertainty properties Margatina et al. (2023).
4.4 Implementation Details
We conduct the example selection exclusively on the training set, ensuring that we remain blind to both the validation and test sets at this stage. After that, we evaluate these candidates on the validation set to determine which uncertainty category performs best and adopt the best-performing category example for ICL testing. All ICL experiments are repeated three times with different random seeds, and we report the average accuracy along with the standard deviation. As for the prompt for evaluating the baseline method, we use the same as the one with no-label setting.
5 Experimental Results and Analysis
First, we conduct experiments with several methods that measure output inconsistency to demonstrate that uncertain instances are more informative than the certain ones when used as in-context examples. Subsequently, we compare our Unc-TTP-guided uncertainty sampling strategy with other common baselines, to further illustrate the validity of our method.
5.1 Preliminary Study: Uncertain Versus Certain Instances
Methods Cat. Llama-2 Mistral GPT-3.5
Verification-Based
P(True) CerP 66.6 (1.1) 77.5 (1.2) 81.4 (1.2)
UncP 67.0 (1.0) 76.4 (0.8) 84.4 (1.3)
SelfCheck CerS 66.7 (1.7) 77.5 (2.4) 83.8 (1.3)
GPT UncS 67.1 (1.6) 76.4 (2.2) 81.9 (1.6)
Sampling-Based
Vanilla CerW 66.7 (1.7) 78.1 (2.9) 84.4 (1.9)
CerR 66.9 (2.0) 78.2 (3.4) 83.7 (2.0)
Sampling Unc 67.6 (1.1) 79.5 (1.2) 85.4 (2.9)
Unc-TTP CerW 69.0 (4.4) 79.5 (3.4) 82.3 (2.3)
CerR 68.1 (2.7) 75.2 (2.2) 84.6 (2.4)
Unc 70.3 (2.2) 81.5 (1.5) 85.2 (1.5)
Table 2: The average accuracy on three datasets for each model over four methods that measure uncertainty via output inconsistency.
Methods Llama-2 Mistral GPT-3.5
SH HS FP Avg. SH HS FP Avg. SH HS FP Avg.
Random Baseline
Random 69.8 (1.6) 53.5 (1.5) 78.3 (1.8) 67.2 (1.6) 71.2 (1.6) 81.6 (4.4) 84.6 (2.3) 79.1 (2.8) 83.1 (4.3) 88.8 (1.8) 83.2 (0.8) 85.0 (2.3)
Traditional Active Learning Methods
Similarity* 69.9 (0.7) 52.7 (1.5) 77.3 (0.0) 66.6 (0.7) 76.0 (0.5) 78.8 (0.6) 86.1 (0.3) 80.3 (0.5) 79.2 (0.6) 87.3 (0.6) 83.4 (0.2) 83.3 (0.5)
BM25* 68.0 (0.7) 55.4 (1.1) 74.7 (1.2) 66.0 (1.0) 72.8 (0.6) 80.0 (1.1) 85.5 (0.9) 79.4 (0.9) 74.8 (1.6) 87.2 (1.0) 83.2 (1.0) 81.7 (1.2)
Diversity 66.5 (1.0) 55.9 (0.7) 81.5 (0.6) 68.0 (0.8) 53.7 (0.9) 81.1 (1.9) 84.1 (0.6) 73.0 (1.1) 62.7 (1.6) 86.7 (0.9) 81.7 (1.1) 77.0 (1.2)
Entropy✝ 68.5 (0.8) 52.7 (0.1) 80.2 (0.9) 67.1 (0.6) 68.0 (0.4) 81.5 (0.7) 83.0 (0.7) 77.5 (0.6) - - - -
Perplexity✝ 66.3 (1.3) 54.9 (1.6) 80.5 (0.7) 67.2 (1.2) 69.0 (0.7) 79.5 (0.3) 85.2 (1.3) 77.9 (0.8) - - - -
Inconsistency-Defined Uncertainty Based Active Learning Methods
P(true) 68.2 (0.8) 51.5 (1.2) 81.2 (1.1) 67.0 (1.0) 73.2 (1.0) 80.7 (0.6) 75.2 (0.8) 76.4 (0.8) 77.8 (1.0) 92.2 (0.6) 83.3 (2.2) 84.4 (1.3)
SelfCheckGPT 66.2 (1.7) 57.3 (1.9) 77.8 (1.1) 67.1 (1.6) 65.5 (2.9) 80.8 (2.1) 82.9 (1.5) 76.4 (2.2) 77.3 (3.5) 87.5 (0.7) 81.0 (0.7) 81.9 (1.6)
Vanilla Sampling 69.1 (1.0) 53.8 (1.0) 79.8 (1.2) 67.6 (1.1) 72.0 (0.9) 79.5 (1.4) 87.0 (1.2) 79.5 (1.2) 80.1 (5.5) 92.5 (1.8) 83.5 (1.3) 85.4 (2.9)
Unc-TTP 71.2 (2.1) 58.8 (3.2) 80.8 (1.3) 70.3 (2.2) 74.7 (0.8) 81.8 (2.9) 88.1 (0.8) 81.5 (1.5) 77.8 (2.4) 93.0 (0.5) 84.8 (1.7) 85.2 (1.5)
Table 3: 1-shot results of LLMs on three subjective tasks in the test set. The table presents the mean accuracy across three estimations, with standard deviations in parentheses. The highest accuracy values under each method are underlined. The best results for each dataset are highlighted in bold. Please note that * indicates methods that select different examples for each test instance, while ✝ signifies methods that require white-box access, which is not compatible with the closed-source LLM GPT-3.5.
To justify the uncertainty defined by inconsistency, we draw on the concept of uncertainty sampling in AL and apply inconsistency-defined uncertainty on in-context example selection to evaluate its effect. The results are shown in Table 2 and visualized in Figure 4.
Verification-based methods try to measure the consistency between the LLM’s first stage answer and the second stage self-verification to reveal LLM uncertainty. In this case, risking more possibility of sycophancy and cannot reflect the intrinsic uncertainty of LLM. Therefore showing no clear pattern of the relative improvement in accuracy of uncertainty examples compared to certainty examples in ICL for each model.
However, sampling-based methods, vanilla sampling and our Unc-TTP, select uncertainty instances which bring consist performance gain compare to the certain ones in most of the cases, showcasing that the selected uncertain instances are indeed more informative than the certain one under the context of ICL. Moreover, compared to vanilla, the relative performance gain is more significant for the uncertain example classified by our proposed Unc-TTP approach, which introduces label injection to strengthen the sampling rigor. As for the exception, SH on GPT-3.5, we will discuss it on the following section.
5.2 Main Results: Comparing Unc-TTP with Previous AL Methods
We have shown that the uncertainty examples selected through our Unc-TTP are more informative than the certainty examples on ICL. In this section we continue to explore the effectiveness of our proposed Unc-TTP compared to other AL strategies. The results are shown in Table 3.
For applying the traditional AL strategy on ICL, Similarity have demonstrated a obvious advantage over other strategy on GPT-3.5 and Mistral, even with the more common strategy like Entropy in traditional AL, which align with the previous results of Margatina et al. (2023). It suggests that the more capable models rely more on sample similarity while weaker model like Llama-2 depends more on example Diversity. Yet, we interestingly discover, compare to other traditional AL strategies, Random are the most informative on the strongest model GPT-3.5. It indicates that strong LLM may be insensitive to examples.
For inconsistency-defined uncertainty as the active in-context example selection strategy, ours Unc-TTP demonstrate a competitive advantage. Moreover, Unc-TTP not only performs better than the AL baseline Random by the average of 3.1%, 2.4% and 0.2% on Llama-2, Mistral and GPT-3.5, but also far exceeds the best strategy Similarity stated in previous work Margatina et al. (2023) with the improvement of 3.7%, 1.2% and 1.9%, demonstrating the effectiveness of our Unc-TTP on selecting the most informative in-context examples by the inconsistency-defined uncertainty.
5.3 Assessing the Scalability of Unc-TTP
To explore the informativeness of the selected inconsistency-defined uncertain in-context examples as the number of shot increases, we conduct $N$-shot experiments on Mistral, taking the AL baseline method Random as comparison.
The results, depicted in Figure 5, reveal that uncertainty examples identified by Unc-TTP are generally more informative than those selected by the vanilla sampling-based method or chosen randomly. Unc-TTP consistently outperforms other baselines across 1 to 4 shots, highlighting the informativeness of the examples selected under Unc-TTP’s guidance. Although the advantages of inconsistency-defined uncertainty sampling diminish for the SH and FP datasets when the number of shots exceeds 8, we argue that this result is reasonable because as the number of shots increases, it is impossible to avoid introducing other variables, such as diversity and length of the instances, and the current exploration of the best combination of in-context examples is still ongoing, which will also be the direction of our future work.
5.4 Assessing the Transferability of Unc-TTP
We have shown that the wavering behavior of LLMs under limited sampling reflects their intrinsic uncertainty, and we leverage this inconsistency-defined uncertainty for active in-context example selection. In Section 3.1, we observed that more capable LLMs better resist skepticism or incorrect information, maintaining their stance while achieving higher accuracy. Based on this, we explore whether Unc-TTP selects more informative uncertain instances—where stronger LLMs waver—that provide stronger guidance for weaker LLMs compared to their self-guided instances.
[…]
Method SH HS FP Average
@@ 5.4 Assessing the Transferability of Unc-TTP
Random 69.8 (1.6) 53.5 (1.5) 78.3 (1.8) 67.2 (1.6)
UncGPT3.5 67.8 (1.5) 58.8 (1.6) 81.2 (2.2) 69.3 (1.8)
UncGPT4 71.0 (2.1) 56.8 (2.6) 80.7 (1.5) 69.5 (2.1)
UncLlama2 71.2 (2.1) 58.8 (3.2) 80.8 (1.3) 70.3 (2.2)
@@
Random 71.2 (1.6) 81.6 (4.4) 84.6 (2.3) 79.1 (2.8)
UncGPT3.5 71.7 (3.6) 79.6 (1.6) 86.9 (2.0) 79.4 (3.3)
UncGPT4 75.0 (2.0) 77.1 (4.9) 85.8 (2.4) 79.3 (3.1)
UncMistral 74.7 (0.8) 81.8 (2.9) 88.1 (0.8) 81.5 (1.5)
Table 4: Transferability of Unc-TTP-guided uncertainty sampling in 1-shot ICL: Comparing the informativeness of uncertainty examples selected by strong models versus those selected through weak model self-guidance.
@@ 6 Conclusion
In this paper, we introduce the Unc-TTP, a novel method for assessing intrinsic uncertainty in both open- and closed-source LLMs. From the perspective of applying active learning principles to in-context example selection, we experimentally demonstrate that uncertainty examples selected through Unc-TTP are more informative than certainty examples. This provides further evidence that output inconsistency can serve as an indicator of LLMs’ intrinsic uncertainty.
Additionally, we propose an uncertainty-based active in-context example selection strategy grounded in Unc-TTP, which outperforms previous AL strategies in terms of performance gains in ICL. These findings establish Unc-TTP as an effective tool for classifying LLM uncertainty and enhancing ICL performance, paving the way for understanding the uncertainty behavior of LLMs.
@@
Firstly, our proposed Unc-TTP inherits the prompt-sensitive nature of the LLMs, which may pose a risk when applying Unc-TTP-guided uncertainty sampling. However, for the robustness of our experiments, we repeated our experiments three times and reported the mean and standard deviation.
Secondly, we have only evaluated Unc-TTP on subjective tasks. Given highly subjective tasks often require a deep understanding of contextual subtleties, their labels are sometimes "biased" themselves depending on the annotator decomposition (Li et al., 2022, 2023). This characteristic is more compatible with Unc-TTP, as it can effectively expose areas where LLMs exhibit uncertainty compared to objective reasoning tasks. For the implementation of Unc-TTP on objective tasks will be a primary focus in our future work.
Thirdly, we have only tested Unc-TTP on classification tasks, as designing intuitive perturbations, like label injection, for generative or more complex tasks is challenging. This is because such perturbations inevitably introduce additional variables. One possible approach is to transform the LLM’s output into a binary fact-checking task (e.g., SelfCheckGPT). However, our paper shows that this verification-based method is ineffective under the scenario of in-context example selection. Furthermore, Unc-TTP is specifically designed for active context example selection, whereas complex generation tasks and ICL may not be compatible. Developing appropriate perturbations to extend Unc-TTP to more general generative tasks will be the focus of our future work.
Finally, although Unc-TTP adopts a one-example-for-all approach, which makes inference more efficient compared to retrieval-based methods like similarity, it requires more resources for sample selection. However, it is important to note that Unc-TTP not only serves as a novel active example selection method but also as an uncertainty classification method that can reflect the LLM’s intrinsic uncertainty. In the future, we aim to explore additional application scenarios for Unc-TTP, a perturbation-based, inconsistency-defined uncertainty classification method.
@@
Appendix A Detailed Performance of LLMs under Preliminary Test
The results of preliminary test under three label injection settings on three benchmarks are shown in Figure 2. LLM shows more compliance with correct labels. For the setting that provides labels, all models except Llama-2 on the SH dataset show gullibility to the given labels. Particularly, in the right-label setting, some models nearly reach 100% agreement with the given label. However, for the wrong-label setting, the model exhibits varying levels of resistance to enticement. Besides, the more capable the model is, the better it maintains accuracy in the wrong-label setting, although it does not converge to the accuracy achieved in the no-label setting. This indicates that unless the LLM is very certain of its answer, it reacts quite prudently when the labels provided deviate considerably from its own answers, as adopting the user’s point of view seems to be a safer option.
LLMs tend to make compromises when conservative options are offered, choosing the neutral option to accommodate the user’s preference. In the FP dataset, all models exhibit relatively steady performance under the wrong-label setting. A possible explanation for this observation is that FP is a classification dataset with three labels, where the distinctions between neutral and positive or negative are subtle. LLMs inherently tend to favor the more conservative neutral answers. When their initial beliefs are challenged in the wrong-label setting, they are inclined to prefer the neutral option, potentially aligning with user preferences. This behavior corresponds to the data distribution in Table 6 of the Appendix. We observe that most LLMs exhibit a tendency to produce conservative answers in the no-label setting. This tendency becomes pronounced in the wrong-label setting, where the number of neutral answers doubles, further substantiating our conclusion.